# ShadowMaskFormer: Mask Augmented Patch Embedding for Shadow Removal

### v2

@@ Abstract
Transformer recently emerged as the de facto model for computer vision tasks and has also been successfully applied to shadow removal. However, these existing methods heavily rely on intricate modifications to the attention mechanisms within the transformer blocks while using a generic patch embedding. As a result, it often leads to complex architectural designs requiring additional computation resources. In this work, we aim to explore the efficacy of incorporating shadow information within the early processing stage. Accordingly, we propose a transformer-based framework with a novel patch embedding that is tailored for shadow removal, dubbed ShadowMaskFormer. Specifically, we present a simple and effective mask-augmented patch embedding to integrate shadow information and promote the model’s emphasis on acquiring knowledge for shadow regions. Extensive experiments conducted on the ISTD, ISTD+, and SRD benchmark datasets demonstrate the efficacy of our method against state-of-the-art approaches while using fewer model parameters.
@@ 1 Introduction
Limitation of Shadow Transformer. Recently, as an emerging backbone model of choice for vision tasks, transformers have also been applied to shadow removal tasks. However, there are two issues. ❶ The existing transformer-based methods overlook the shadow information in the early processing stage and directly employ the generic patch embedding, which cannot fully unleash the representational power of the transformer-based methods. As shown in Figure 1(a), the state-of-the-art transformer-based methods (i.e., CRFormer [12] and ShadowFormer [13]) achieve the contextual knowledge from the non-shadow regions, resulting in performance degradation. ❷ As shown in Figure 1(c)(c), the transformer-based methods mainly necessitate the creation of new modules with shadow masks within the main computation blocks (i.e., transformer blocks and convolution blocks) for shadow removal, resulting in a noticeable scale of parameters. Hence, this has led to the following research question: Can we incorporate shadow information in patch embedding to avoid complicated modifications to transformer blocks and highlight the shadow region in the feature extraction stage?
@@
ShadowMaskFormer. To overcome the constraints of current transformer-based techniques and address the research question, we introduce a new framework called ShadowMaskFormer. As shown in Figure 1(b), ShadowMaskFormer combines the transformer model with a shadow mask in patch embedding, presenting a novel approach to effectively remove shadows from images. Specifically, in the early processing stage (i.e., patch embedding), we propose a simple and effective patch embedding, namely Mask Augmented Patch Embedding (MAPE). The motivation behind MAPE is our observations based on the limited utilization of masks from existing work, as detailed in Section 3.2. Therefore, in our ShadowMaskFormer, the shadow mask is carefully utilized in MAPE with two complementary binarization schemes (the and Binarization) to enhance the shadow region pixels, as detailed in section 4.2. In conclusion, with the proposed MAPE, our ShadowMaskFormer can not only leverage the contextual information that is achieved by the learning capability of the transformer model but also restore the shadow region pixels at an earlier time. This approach allows the model to more effectively acquire the distinctive features of shadow regions and perform shadow removal with more purposefully. It is noteworthy that MAPE operates as a single-layer module, primarily performing pixel-level operations, and it requires computation only once per training epoch, significantly reducing computational demands. As far as our knowledge extends, we believe our approach marks the pioneering exploration of utilizing patch embedding in vision transformer models for the task of shadow removal. Experimental results demonstrate that ShadowMaskFormer achieves outstanding shadow removal results over the three widely-used shadow removal datasets, surpassing the state-of-the-art performance. Furthermore, compared with other SOTA methods, it only has 2.2MB network parameters, e.g., half of CRFormer [12] and smaller than others, as illustrated in Figure 1(c). The main contributions of this work are as follows:
We propose an innovative transformer-based framework called ShadowMaskFormer111Our implementation is available at https://www.github.com/lizhh268/ShadowMaskFormer.git, which incorporates mask information in the patch embedding stage. To the best of our knowledge, we are the first to explore the task of shadow removal from the perspective of patch embedding in vision transformer models.
We propose the concept patch embedding tailored for shadow removal and further introduce the corresponding MAPE which is simple and effective. By carefully utilizing the shadow mask with two complementary binarization schemes, MAPE effectively enhances the shadow region to assist in restoring each pixel without introducing any model parameters.
Comprehensive experiments conducted on publicly available datasets, namely ISTD, ISTD+, and SRD, demonstrate the outstanding performance of our proposed method, placing it at the state-of-the-art level.
@@ 2.1 Shadow Removal
In the recent past, deep learning-based (DL) methods have achieved excellent performance in the field of shadow removal, leveraging their end-to-end capabilities.
DL methods can be primarily categorized into two main approaches: convolutional neural network (CNN) based and generative modeling, particularly generative adversarial network (GAN), based approaches. CNN-based approaches employ multi-level convolutions to extract context-based features for shadow removal [5, 7, 21, 22, 23, 24, 6, 25, 26]. For instance, DeshadowNet [5] utilizes image contextual information to learn shadow mask features and remove shadows. GAN-based approaches generate shadow mask images and shadow-free images that closely resemble reality by adhering to a series of criteria for discerning between real and fake data [9, 11, 8, 10, 27, 28, 29, 28]. For instance, in DC-ShadowNet [28], three dedicated losses were defined to characterize shadow images for accurate shadow synthesis. Besides, the most recent work [30] predicts the adaptive weights between two features that are extracted from the shadow image and the shadow mask image. Our method is different from them, i.e., we define the problem as the integration of transformer models and shadow masks, which enables us to draw inspiration from existing mask utilization methods.
@@ 2.3 Transformer for Shadow Removal
With the remarkable performance of transformer models in computer vision tasks, it has also been applied to shadow removal tasks [12, 13]. CRFormer [12] attempts to guide the restoration of shadow regions by leveraging non-shadow region information with shadow masks. However, it places excessive emphasis on non-shadow region information, resulting in insufficient model attention allocation to the crucial shadow regions and shadow masks. ShadowFormer [13] proposes a Shadow-Interaction Module Attention to exploit the global contextual correlation between shadow and non-shadow regions. It is worth noting that the transformer-based method, still like other DL-based methods, relies on designing new modules from the transformer blocks to leverage shadow masks for the shadow removal task, as shown in Figure 1(a). In this work, we aim to explore the efficacy of incorporating shadow information during the early processing stage and propose a novel patch embedding module tailored to the shadow removal task.
@@ 3.1 Background
In conventional computer vision models, such as Convolutional Neural Networks (CNNs), an input image is processed as an entirety by sliding a filter over the entire image to capture local patterns. In contrast, a vision transformer (ViT) model [32] does not directly operate on an input image, rather, it breaks down an input image into smaller and non-overlapping patches. Each patch is then linearly transformed into a fixed-dimensional vector representation for subsequent computation. This process is known as Patch Embedding.
@@ 3.2 Motivation
Shadow masks. With a steady stream of promising empirical results confirming the effectiveness of shadow masks for shadow removal [8, 11, 12, 13], we observe two patterns emerged from existing works: ❶ Shadow masks are typically represented as binary masks (i.e., with being the size of an input image) where “” indicates shadow regions and “” indicates non-shadow regions, respectively [8, 23, 12, 13]; ❷ Shadow masks are primarily utilized by the main computing units of the models, e.g., convolution operations in CNNs [40, 23] and attention modules in ViTs [12, 13].
@@
In the above expression, the gain factor in our method is derived from the assumption underlying Eq. 1, and will be learned by the model for the shadowed pixel . Specifically, considering the natural attribute of shadows where the attenuation of light sources results in shadow pixel values significantly lower than those of non-shadowed pixels, we assume that in Eq. 1, will be noticeably greater than 1, and will be smaller compared to . The term appears to be roughly equivalent to the body reflection. Hence, the crux of shadow removal lies in ensuring that the gain factor learned by the model closely approximates the correct solution. In this paper, we endeavor to perform a preliminary exploration of the gain factor during the patch embedding stage. This approach allows the model to enhance shadow removal performance more effectively at an early stage, thereby reducing unnecessary model exploration. Specifically, to address the task of shadow removal, we introduce a simple and effective patch embedding, namely Mask Augmented Patch Embedding (MAPE).
@@ 4 ShadowMaskFormer
In this section, we will elucidate the overarching workflow framework of the model, with a particular emphasis on detailing the implementation of MAPE.
@@ 4.2 Mask Augmented Patch Embedding
Based on the discussions above, aiming to find an accurate approximation of the gain factor for shadow region pixels early in the model’s training, we introduce the concept of Mask Augment Patch Embedding (MAPE) which is based on the two complementary binarization schemes (the and Binarization) as depicted in Figure 2. In particular, to ensure the learned non-linear mapping function of the model to explore the correct early and avoid exploring useless information for shadow region enhancement, we designed MAPE to be positioned in the early stage of the model. In the patch embedding stage, the shadow image and its corresponding shadow mask M are first used as inputs to the model. Subsequently, the and (Binarize) pixel-wise operations are performed on M, resulting in two complementary masks and , respectively. These operations can be expressed using the following formulas:
@@ 5.1 Experimental Setup
Baselines. Towards obtaining a representative evaluation of our method, we consider a wide range of baselines, including a traditional method [45], generative modeling-assisted approaches [11, 8, 29, 10, 9, 30], two transformer-based approaches [12, 13], among others [7, 28, 22].
Evaluation metrics. We use root mean square error (RMSE) between the estimated shadow-free images and the ground truth in the LAB color space, where lower values indicate better results. We also use the peak signal-to-noise ratio (PSNR) and the structural similarity (SSIM) [46] metrics to quantitatively compare the performance of various methods in the RGB color space, where higher values indicate better results.
@@
Method Params Shadow Region (S) Non-Shadow Region (NS) All Image (All) (M) PSNR SSIM RMSE PSNR SSIM RMSE PSNR SSIM RMSE Input Image - 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05 DeshadowNet [5] - - - 11.78 - - 4.84 - - 6.64 DSC [7] 22.3 30.65 0.960 8.62 31.94 0.956 4.41 27.76 0.903 5.71 DHAN [10] 21.8 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67 Fu et al. [6] 187 32.26 0.966 9.55 31.87 0.945 5.74 28.40 0.893 6.50 DC-ShadowNet [28] 21.2 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65 CANet [22] 358 - - 7.82 - - 5.88 - - 5.98 SG-shadowNet [24] 6.20 - - 7.53 - - 2.97 - - 4.23 Zhu et al. [23] 10.1 34.94 0.980 7.44 35.85 0.982 3.74 31.72 0.952 4.79 CRFormer [12] 4.89 - - 7.14 - - 3.15 - - 4.25 ShadowFormer [13] 2.40 36.13 0.988 6.05 35.95 0.986 3.55 32.38 0.955 4.09 MAPE (ours) 2.28 37.71 0.988 5.55 38.23 0.984 2.98 34.43 0.968 3.64
@@ 5.2 Experimental Results
Results on ISTD. Table 1 summarizes the results on the ISTD dataset. In general, we observe that our method consistently outperforms all considered baselines on different regions (i.e., shadow and non-shadow regions) under both input resolution settings (i.e., and ) with fewer model parameters. In particular, our method achieves 1.6 lower RMSE (averaged over whole images) with 9.3 fewer parameters than DC-ShadowNet [28]. Our method also achieves 1.4 lower RMSE (averaged over whole images) with 2.1 fewer parameters than CRFormer [12]. Our method is also all-around competitive against ShadowFormer [13].
@@
Results on ISTD+. Table 2 summarizes the results on the ISTD+ dataset. In general, we observe that (i) all methods exhibit better performance than that on the original ISTD dataset [8] after adjustments of illumination inconsistency introduced by [21]; (ii) our method continues to outperform all the considered baseline competitors as indicated by RMSE while using fewer model parameters. Specifically, our method achieves better RMSE on shadow, non-shadow, and all regions than both SG-ShadowNet [24] and ShadowFormer [13].
Results on SRD. Table 3 presents the comparison results on the SRD [5] dataset. Evidently, we observe that our method achieves significantly better results on the overall image than all considered baselines. For instance, our method yields 1.58dB to 2.28dB PSNR improvements on shadow, non-shadow, and all regions over ShadowFormer [13] with a similar number of model parameters. In addition, we observe that our method also achieves noticeably better RMSE than the baselines. Specifically, our method achieves 1.3 and 1.2 lower RMSE than CRFormer [12] on shadow and all regions, respectively, while using 2.1 fewer model parameters.

### v3

@@ Abstract
Transformer recently emerged as the de facto model for computer vision tasks and has also been successfully applied to shadow removal. However, these existing methods heavily rely on intricate modifications to the attention mechanisms within the transformer blocks while using a generic patch embedding. As a result, it often leads to complex architectural designs requiring additional computation resources. In this work, we aim to explore the efficacy of incorporating shadow information within the early processing stage. Accordingly, we propose a transformer-based framework with a novel patch embedding that is tailored for shadow removal, dubbed ShadowMaskFormer. Specifically, we present a simple and effective mask-augmented patch embedding to integrate shadow information and promote the model’s emphasis on acquiring knowledge for shadow regions. Extensive experiments conducted on the ISTD, ISTD+, and SRD benchmark datasets demonstrate the efficacy of our method against state-of-the-art approaches while using fewer model parameters. Our implementation is available at https://github.com/lizhh268/ShadowMaskFormer.
@@ 1 Introduction
Limitation of Shadow Transformer. Recently, as an emerging backbone model of choice for vision tasks, transformers have also been applied to shadow removal tasks. However, there are two issues. ❶ The existing transformer-based methods overlook the shadow information in the early processing stage and directly employ the generic patch embedding, which cannot fully unleash the representational power of the transformer-based methods. As shown in Figure 1(a), the state-of-the-art transformer-based methods (i.e., CRFormer [12] and ShadowFormer [13]) achieve the contextual knowledge from the non-shadow regions, resulting in performance degradation. ❷ As shown in Figure 1(c), the transformer-based methods mainly necessitate the creation of new modules with shadow masks within the main computation blocks (i.e., transformer blocks and convolution blocks) for shadow removal, resulting in a noticeable scale of parameters. Hence, this has led to the following research question: Can we incorporate shadow information in patch embedding to avoid complicated modifications to transformer blocks and highlight the shadow region in the feature extraction stage?
@@
ShadowMaskFormer. To overcome the constraints of current transformer-based techniques and address the research question, we introduce a new framework called ShadowMaskFormer. As shown in Figure 1(b), ShadowMaskFormer combines the transformer model with a shadow mask in patch embedding, presenting a novel approach to effectively remove shadows from images. Specifically, in the early processing stage (i.e., patch embedding), we propose a simple and effective patch embedding, namely Mask Augmented Patch Embedding (MAPE). The motivation behind MAPE is our observations based on the limited utilization of masks from existing work, as detailed in Section 3.2. Therefore, in our ShadowMaskFormer, the shadow mask is carefully utilized in MAPE with two complementary binarization schemes (the and Binarization) to enhance the shadow region pixels, as detailed in section 4.2. In conclusion, with the proposed MAPE, our ShadowMaskFormer can not only leverage the contextual information that is achieved by the learning capability of the transformer model but also restore the shadow region pixels at an earlier time. This approach allows the model to more effectively acquire the distinctive features of shadow regions and perform shadow removal more purposefully. It is noteworthy that MAPE operates as a single-layer module, primarily performing pixel-level operations, and it requires computation only once per training epoch, significantly reducing computational demands. As far as our knowledge extends, we believe our approach marks the pioneering exploration of utilizing patch embedding in vision transformer models for the task of shadow removal. Experimental results demonstrate that ShadowMaskFormer achieves outstanding shadow removal results over the three widely-used shadow removal datasets, surpassing the state-of-the-art performance. Furthermore, compared with other SOTA methods, it only has 2.2MB network parameters, e.g., half of CRFormer [12] and smaller than others, as illustrated in Figure 1. The main contributions of this work are as follows:
We propose ShadowMaskFormer, a novel framework that integrates shadow mask information in the patch embedding stage. To the best of our knowledge, this is the first attempt to incorporate shadow-specific information at such an early stage in transformer models, establishing a new paradigm for shadow removal tasks.
We present MAPE, a simple yet effective method tailored for integrating shadow information, naturally aligned with the patch embedding stage. By leveraging shadow masks with two complementary binarization schemes, MAPE enhances shadow region representations, enabling precise pixel-wise restoration without adding extra parameters. This results in both computational efficiency and robust performance across a wide range of scenarios.
Our method achieves state-of-the-art performance across benchmark datasets (ISTD, ISTD+, and SRD), demonstrating both exceptional accuracy and computational efficiency with a model size of only 2.2MB. Extensive experiments further demonstrate the robustness of MAPE even under incomplete or inaccurate shadow masks, proving its outstanding ability to generalize to complex, real-world conditions.
@@ 2.1 Shadow Removal
In the recent past, deep learning-based methods have achieved excellent performance in the field of shadow removal, leveraging their end-to-end capabilities.
Deep learning-based methods can be primarily categorized into two main approaches: convolutional neural network (CNN) based and generative modeling, particularly generative adversarial network (GAN), based approaches. CNN-based approaches employ multi-level convolutions to extract context-based features for shadow removal [5, 7, 21, 22, 23, 24, 6, 25, 26]. For instance, DeshadowNet [5] utilizes image contextual information to learn shadow mask features and remove shadows. GAN-based approaches generate shadow mask images and shadow-free images that closely resemble reality by adhering to a series of criteria for discerning between real and fake data [9, 11, 8, 10, 27, 28, 29, 28]. For instance, in DC-ShadowNet [28], three dedicated losses were defined to characterize shadow images for accurate shadow synthesis. Besides, AdaptiveFusionNetwork [30] predicts the adaptive weights between two features that are extracted from the shadow image and the shadow mask image. On the other hand, two recent works have made promising progress in shadow removal: ShadowDiffusion [31] achieves notable performance by proposing a dynamic mask-aware diffusion model and incorporating degradation priors, while HomoFormer [32] offers another solution by using random shuffle and reverse shuffle operations to transform non-uniform shadows into a uniform distribution, enhancing the effectiveness of local self-attention for complex shadows. Our method is different from them, i.e., we define the problem as the integration of transformer models and shadow masks, which enables us to draw inspiration from existing mask utilization methods.
@@ 2.3 Transformer for Shadow Removal
With the remarkable performance of transformer models in computer vision tasks, it has also been applied to shadow removal tasks [12, 13]. CRFormer [12] attempts to guide the restoration of shadow regions by leveraging non-shadow region information with shadow masks. However, it places excessive emphasis on non-shadow region information, resulting in insufficient model attention allocation to the crucial shadow regions and shadow masks. ShadowFormer [13] proposes a Shadow-Interaction Module Attention to exploit the global contextual correlation between shadow and non-shadow regions. It is worth noting that the transformer-based method, still like other deep learning-based methods, relies on designing new modules from the transformer blocks to leverage shadow masks for the shadow removal task, as shown in Figure 1(a). In this work, we aim to explore the efficacy of incorporating shadow information during the early processing stage and propose a novel patch embedding module tailored to the shadow removal task.
@@ 3.1 Background
In conventional computer vision models, such as convolutional neural networks, an input image is processed as an entirety by sliding a filter over the entire image to capture local patterns. In contrast, a vision transformer (ViT) model [34] does not directly operate on an input image, rather, it breaks down an input image into smaller and non-overlapping patches. Each patch is then linearly transformed into a fixed-dimensional vector representation for subsequent computation. This process is known as Patch Embedding.
@@ 3.2 Motivation
Shadow masks. With a steady stream of promising empirical results confirming the effectiveness of shadow masks for shadow removal [8, 11, 12, 13], we observe two patterns emerged from existing works: ❶ Shadow masks are typically represented as binary masks (i.e., with being the size of an input image) where “” indicates shadow regions and “” indicates non-shadow regions, respectively [8, 23, 12, 13]; ❷ Shadow masks are primarily utilized by the main computing units of the models, e.g., convolution operations in CNN-based models [42, 23] and attention modules in ViT models [12, 13].
@@
In the above expression, the gain factor in our method is derived from the assumption underlying Eq. 1, and will be learned by the model for the shadowed pixel . Specifically, considering the natural attribute of shadows where the attenuation of light sources results in shadow pixel values significantly lower than those of non-shadowed pixels, we assume that in Eq. 1, will be noticeably greater than 1, and will be smaller compared to . The term appears to be roughly equivalent to the body reflection. Hence, the crux of shadow removal lies in ensuring that the gain factor learned by the model closely approximates the correct solution. In this paper, we endeavor to perform a preliminary exploration of the gain factor during the patch embedding stage. This approach allows the model to enhance shadow removal performance more effectively at an early stage, thereby reducing unnecessary model exploration. Specifically, to address the task of shadow removal, we introduce a simple and effective patch embedding, namely Mask Augmented Patch Embedding.
@@ 4 ShadowMaskFormer
In this section, we will elucidate the overarching workflow framework of the model, with a particular emphasis on detailing the implementation of Mask Augmented Patch Embedding.
@@ 4.2 Mask Augmented Patch Embedding
Based on the discussions above, aiming to find an accurate approximation of the gain factor for shadow region pixels early in the model’s training, we introduce the concept of Mask Augmented Patch Embedding which is based on the two complementary binarization schemes (the and Binarization) as depicted in Figure 2. In particular, to ensure the learned non-linear mapping function of the model to explore the correct early and avoid exploring useless information for shadow region enhancement, we designed MAPE to be positioned in the early stage of the model. In the patch embedding stage, the shadow image and its corresponding shadow mask M are first used as inputs to the model. Subsequently, the and (Binarize) pixel-wise operations are performed on M, resulting in two complementary masks and , respectively. These operations can be expressed using the following formulas:
@@ 4.3 Loss Function
Following prior shadow removal works [21, 24, 12], we only adopt the image consistency loss , which is mathematically defined as follows:
where and are the predicted shadow-free image and the ground truth, respectively. The use of loss is motivated by its ability to effectively capture pixel-wise differences, which is crucial for accurately removing shadows while preserving fine details. This makes loss well-suited to meet the requirements of our proposed framework.
@@ 5.1 Experimental Setup
Baselines. Towards obtaining a representative evaluation of our method, we consider a wide range of baselines, including a traditional method [47], generative modeling-assisted approaches [11, 8, 29, 10, 9, 30], two transformer-based approaches [12, 13], among others [7, 28, 22, 31].
Evaluation metrics. Following the previous methods [21, 25, 32], we use mean absolute error (MAE) between the estimated shadow-free images and the ground truth in the LAB color space, where lower values indicate better results. We also use the peak signal-to-noise ratio (PSNR) and the structural similarity (SSIM) [48] metrics to quantitatively compare the performance of various methods in the RGB color space, where higher values indicate better results.
@@
Method Params Shadow Region (S) Non-Shadow Region (NS) All Image (All) (M) PSNR SSIM MAE PSNR SSIM MAE PSNR SSIM MAE Input Image - 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05 DeshadowNet [5] - - - 11.78 - - 4.84 - - 6.64 DSC [7] 22.3 30.65 0.960 8.62 31.94 0.956 4.41 27.76 0.903 5.71 DHAN [10] 21.8 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67 Fu et al. [6] 187 32.26 0.966 9.55 31.87 0.945 5.74 28.40 0.893 6.50 DC-ShadowNet [28] 21.2 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65 CANet [22] 358 - - 7.82 - - 5.88 - - 5.98 SG-shadowNet [24] 6.20 - - 7.53 - - 2.99 - - 4.23 Zhu et al. [23] 10.1 34.94 0.980 7.44 35.85 0.982 3.74 31.72 0.952 4.79 CRFormer [12] 4.89 - - 7.14 - - 3.15 - - 4.25 ShadowDiffusion [31] 55.52 38.72 0.987 4.98 37.78 0.985 3.44 34.73 0.970 3.63 ShadowFormer [13] 2.40 36.13 0.988 6.05 35.95 0.986 3.55 32.38 0.955 4.09 ShadowMaskFormer (ours) 2.28 37.71 0.988 5.55 38.23 0.984 2.98 34.43 0.968 3.64
@@ 5.2 Experimental Results
Results on ISTD. Table 1 summarizes the results on the ISTD dataset. In general, we observe that our method consistently outperforms all considered baselines on different regions (i.e., shadow and non-shadow regions) under both input resolution settings (i.e., and ) with fewer model parameters. In particular, our method achieves 1.6 lower MAE (averaged over whole images) with 9.3 fewer parameters than DC-ShadowNet [28]. Our method also achieves 1.4 lower MAE (averaged over whole images) with 2.1 fewer parameters than CRFormer [12]. Our method is also all-around competitive against ShadowFormer [13].
@@
Results on ISTD+. Table 2 summarizes the results on the ISTD+ dataset. In general, we observe that (i) all methods exhibit better performance than that on the original ISTD dataset [8] after adjustments of illumination inconsistency introduced by [21]; (ii) our method continues to outperform all the considered baseline competitors as indicated by MAE while using fewer model parameters. Specifically, our method achieves better MAE on shadow, non-shadow, and all regions than both SG-ShadowNet [24] and ShadowFormer [13].
Results on SRD. Table 3 presents the comparison results on the SRD [5] dataset. Evidently, we observe that our method achieves significantly better results on the overall image than all considered baselines. For instance, our method yields 1.58dB to 2.28dB PSNR improvements on shadow, non-shadow, and all regions over ShadowFormer [13] with a similar number of model parameters. In addition, we observe that our method also achieves noticeably better MAE than the baselines. Specifically, our method achieves 1.3 and 1.2 lower MAE than CRFormer [12] on shadow and all regions, respectively, while using 2.1 fewer model parameters.